### v1

# RealRep: Generalized SDR-to-HDR Conversion with Style Disentangled Representation Learning

Abstract.
High-Dynamic-Range Wide-Color-Gamut (HDR-WCG) technology is becoming increasingly prevalent, intensifying the demand for converting Standard Dynamic Range (SDR) content to High Dynamic Range (HDR). Existing methods primarily rely on fixed tone mapping operators, which are inadequate for handling SDR inputs with diverse styles commonly found in real-world scenarios. To address this challenge, we propose a generalized SDR-to-HDR method that tackles diverse styles in real-world SDR content, termed Realistic Style Disentangled Representation Learning (RealRep). By disentangling luminance and chrominance, we analyze the intrinsic differences between contents with varying styles and propose a disentangled multi-view style representation learning method. This approach captures the guidance prior of true luminance and chrominance distributions across different styles, even when the SDR style distributions exhibit significant variations, establishing a robust embedding space for inverse tone mapping. Motivated by the difficulty of directly exploiting degradation representation priors, we propose the Degradation-Domain Aware Controlled Mapping Network (DDACMNet)—a two-stage framework that performs adaptive hierarchical mapping guided by a control-aware normalization mechanism. DDACMNet dynamically modulates the mapping process via degradation-conditioned hierarchical features, enabling robust adaptation across diverse degradation domains. Extensive experiments show that RealRep, consistently outperforms state-of-the-art methods with superior generalization and perceptually faithful HDR color gamut reconstruction.
High dynamic range, Wide Color Gamut, inverse tone mapping, Deep learning.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologiesccs: Computing methodologies Reconstructionccs: Computing methodologies Computational photography
Figure 1. An illustration of our motivation. (a) Comparison between previous frameworks (top) and our realistic attributes guidance prior approach (bottom), where the prior incorporates realistic luminance and chrominance guidance to direct the mapping network. (b) t-SNE visualization of the distributions of SDR input features (left) and latent features (right) after fine-tuning the previous methods’ encoder using a multi-degradation dataset.
1. Introduction
Figure 2. Illustration of the superiority of our approach in recovering luminance and chrominance under unknown degradations. (a) Input SDR image, (b) Results of LSNet (Guo et al., 2023) with single degradation training, where the colors closely resemble the input SDR, (c) Results of ICTCPNet (Huang et al., 2023) with multi-degradation training, which still fail to recover realistic brightness and color, (d) Our proposed method (RealRep) generates the best results, with high-quality recovery in both luminance and chrominance, (e) Ground truth HDR image.
High Dynamic Range/Wide Color Gamut (HDR/WCG) media, as defined in (Series, 2012; Standard, 2014), provides an expanded luminance range and a broader spectrum of visible colors, surpassing the limitations of Standard Dynamic Range (SDR) media. With advancements in display technologies, HDR content has shown significant potential in enhancing visual experiences. In the media and film industries, innovations such as advanced Electro-Optical Transfer Functions (OETFs) like PQ and HLG ((2018), ITU), and Wide Color Gamut (WCG) RGB primaries defined by BT.2020 ((2015), ITU), have further amplified the benefits of HDR. However, the vast majority of existing media content remains in SDR format, restricting its potential on HDR-capable displays. Consequently, Inverse Tone Mapping (iTM), which converts SDR to HDR, has become crucial for leveraging HDR display technology and expanding the usability of SDR content.
Initially, iTM approaches relied on rule-based operations (Kim and Kim, 2018; Kim et al., 2019a, b), enabling limited luminance expansion and color enhancement under specific conditions. However, these methods often suffered from artifacts and poor detail reconstruction, particularly in overexposed or underexposed regions. With the advent of deep learning, learning-based methods have emerged as the dominant approach (Chen et al., 2021d; Shao et al., 2022; Xu et al., 2022; Guo et al., 2023), reframing iTM as a mapping problem between SDR and HDR/WCG. These methods have demonstrated significant advancements in visual quality, addressing limitations of earlier approaches.
In previous approaches, priors have been commonly used to address the ill-posed nature of the inverse tone mapping (iTM) task. These priors were typically predefined, relying on fixed Tone Mapping (TM) operators that perform well under specific conditions but fail to generalize across various types of SDR content.Furthermore, many learning-based methods are implicitly trained on specific tone-mapping operators, which limits their generalization capability. For example, tone-mapping styles tailored for online video platforms such as YouTube often prioritize visual attractiveness, whereas degradation from classical operators like Reinhard aims to preserve the visual dynamic range for content production purposes. Consequently, models trained on a single degradation dataset face inherent limitations in robustness and adaptability. This is illustrated in Figure 1(a), where prior methods employ fixed priors that work when paired with predefined TM operators. However, SDR inputs from different sources exhibit significant variations in brightness and color levels, leading to substantial discrepancies in the outputs. Even when the encoder is fine-tuned using a multi-degradation dataset, SDR inputs with highly diverse styles tend to cluster into neighboring groups rather than forming well-separated clusters, as shown in Figure 1(b). This suggests that the prior sampling mechanism in earlier methods is unreliable when applied to real-world SDR inputs with varying degradation types.
To address the aforementioned challenges, we propose RealRep, a novel framework for SDR-to-HDR conversion tailored to real-world scenarios involving complex and diverse degradation types. As shown in Figure 4, SDR degradations from different sources exhibit highly varied trends in brightness and chrominance, highlighting the necessity of constructing a robust disentangled representation space that can accurately capture the intrinsic attribute distributions across styles. To this end, RealRep adopts a style-disentangled representation learning strategy that explicitly separates luminance and chrominance features under diverse degradations, enabling the extraction of realistic and generalizable feature-guided representations for inverse tone mapping (iTM). To further improve representation robustness, we introduce the Degradation-Domain Aware Controlled Mapping Network (DDACMNet). DDACMNet is designed to stably leverage the representation priors learned by RealRep, guiding the model to perform consistent and effective mappings under varying degradation conditions. Moreover, by learning from a wide spectrum of degradation types, RealRep significantly expands its learned representation space to encompass a broader and more perceptually accurate color gamut. This allows the framework to model both large-scale luminance variations and fine-grained chromatic structures, enabling reliable luminance expansion and color restoration in highly unconstrained multi-degradation scenarios.
Our main contributions are summarized as follows:
  • Style disentangled representation learning: We propose RealRep, a novel method that disentangles luminance and chrominance distributions, eliminating the interference caused by differences in diverse styles, and constructs a robust embedding space capable of representing true luminance and chrominance distributions for the SDR-to-HDR task.
  • We introduce the Degradation-Domain Aware Controlled Mapping Network (DDACMNet), a dedicated module designed to enhance mapping consistency under varying degradations. By stably leveraging the representation priors learned by RealRep, DDACMNet facilitates dynamic luminance range expansion and color gamut extension through a global-to-local mapping strategy.
  • Extensive evaluation on diverse real-world SDR styles: Comprehensive experiments on datasets containing both diverse and unknown SDR styles demonstrate that RealRep consistently outperforms state-of-the-art methods, achieving superior visual quality and strong generalization.
Figure 3. Overview of the proposed multi-degradation SDR-to-HDR framework. The architecture consists of a pair of multi-view degradation encoders that extract global and local representations of luminance and chrominance, a multi-view fusion module that facilitates spatial and channel-wise interaction, and the Degradation-Domain Aware Controlled Mapping Network (DDACMNet) comprising degradation-aware mapping and detail refinement. The encoders are jointly trained with a disentanglement objective to ensure cross-domain consistency. The overall framework enables dynamic luminance expansion and perceptually aligned color gamut reconstruction under diverse degradations.
Figure 4. Luminance (a) and chrominance (b) distribution variations across different SDR degradations in the HDRTV4K dataset. Significant distribution differences highlight the need for luminance-chrominance disentangling, enabling more effective representation learning for SDR-to-HDR conversion under diverse degradations.
2. Related Work
2.1. Multiple Styles SDR-to-HDR
Although SDR-to-HDR conversion algorithms for media content and consumer-grade displays hold substantial practical value, they have only recently gained attention and emerge. The earliest work in this area, presented by Kim et al. (Kim and Kim, 2018), addressed the joint execution of super-resolution (SR) and SDR-to-HDR conversion, utilizing convolutional neural networks (CNNs). Subsequently, Deep SR-ITM (Kim et al., 2019a) introduced a modulation block for spatially adaptive local contrast enhancement. JSI-GAN (Kim et al., 2019b) incorporated pixel-level dynamic convolution for contrast modulation and leveraged generative adversarial networks (GANs) to improve visual quality. Chen et al. (Chen et al., 2021d) proposed the HDRTV1K dataset and HDRTVNet, a network combining adaptive global color mapping, local enhancement, and highlight generation. Xu et al. (Xu et al., 2022) introduced FMNet, which integrates DCT-based frequency-aware modulation blocks to address artifacts effectively. More recently, Guo et al. (Guo et al., 2023) proposed the HDRTV4K dataset to address the limitations of existing datasets and introduced LSNet, a luminance segmentation-based solution. They highlighted that existing methods are restricted by their reliance on fixed tone mapping (TM) operators, limiting their ability to generalize across degradations. In this work, we address this limitation through multi-angle representation learning, which extracts generalizable degradation features and enables the network to support multiple SDR-to-HDR mappings without dependence on predefined TM operators.
2.2. Image Enhancement for Multi-Degradations
Image enhancement covers tasks such as super-resolution, denoising, dehazing, deraining, and deblurring. Early deep learning methods typically addressed these tasks separately. For example, SRCNN (Dong et al., 2014) introduced CNNs for super-resolution, followed by deeper and residual-based architectures (Kim et al., 2016; Ahn et al., 2018; Lim et al., 2017). For other tasks, deep models have also shown strong performance in denoising (Tian et al., 2020; Guo et al., 2019; Xixi et al., 2019), dehazing (Li et al., 2017; Zhang and Patel, 2018; Yang et al., 2022), deraining (Hu et al., 2019; Yang et al., 2019; Li et al., 2021), and deblurring (Lu et al., 2019; Eboli et al., 2020; Chen et al., 2020b). To overcome the limitations of task-specific models, recent works have proposed unified frameworks for multi-degradation enhancement. IPT (Chen et al., 2021b), AirNet (Li et al., 2022), and PromptIR (Potlapalli et al., 2023) adopt transformers or prompt-guided strategies to handle diverse degradations within a single model. RAM (Qin et al., 2024) offers a novel perspective by extracting shared content from degraded images directly, without requiring explicit degradation modeling. These unified methods improve generalization and scalability in real-world scenarios with unknown or mixed degradations.
2.3. Degradation Representation
Degradation representation is critical for guiding effective enhancement, especially under diverse or unknown degradation conditions. Traditional methods often assume known degradation types and estimate parameters, such as noise level in denoising (Guo et al., 2018; Liu et al., 2013) or blur kernel in deblurring (Chen et al., 2021c; Gong et al., 2016; Hu et al., 2014). Recent approaches learn implicit representations that encode degradation characteristics without explicit labels. AirNet (Li et al., 2022) leverages contrastive learning, while PromptIR (Potlapalli et al., 2023) conditions restoration with learned prompts. We adopt a similar strategy for multi-style SDR-to-HDR enhancement, learning a degradation-aware representation that captures style-specific information to guide adaptive HDR reconstruction.
3. Methodology
We propose RealRep, a dual-stage framework for SDR-to-HDR restoration, comprising: (1) a representation-learnable multi-view encoder, and (2) a Degradation-Domain Aware Controlled Mapping Network (DDACMNet).
Given an input SDR image $x$, the encoder extracts hierarchical luminance and chrominance features through two modality-specific branches, yielding $e_{\text{lum}}^{g},e_{\text{lum}}^{l}$ and $e_{\text{chr}}^{g},e_{\text{chr}}^{l}$, where the superscripts $g$ and $l$ denote global and local branches, respectively. These features are supervised via contrastive representation learning, as shown in Fig. 3(b), where global and local projectors generate key-anchor pairs $z_{\text{key}}^{g,l},z_{\text{anchor}}^{g,l}$ to ensure semantic consistency across diverse SDR styles. The multi-view encoder outputs are fused using a lightweight Multi-view Fusion Block ($f_{\text{fusion}}$, Fig. 3(c)), producing a unified latent representation $z$. This representation, along with the input image $x$, is processed by DDACMNet ($f_{\text{DDACM}}$), which performs progressive SDR-to-HDR mapping under the control of hierarchical degradation priors $z_{\text{deg}}^{g},z_{\text{deg}}^{l}$. These priors are extracted from encoder features and injected through controlled affine modulation to enable dynamic, degradation-aware restoration. DDACMNet further enhances spatial details and textures through integrated refinement, yielding high-fidelity HDR reconstruction.
The overall process is formulated as:
(1) $\hat{y}=f_{\text{DDACM}}\left(f_{\text{fusion}}\left(e_{\text{lum}}^{g,l},e_{\text{chr}}^{g,l}\right),\;x\right).$
3.1. Realistic Style Disentangled Representation Learning
3.1.1. Multi-View Representation Extraction
To address the diverse SDR degradations, we design a hierarchical multi-view encoder, based on the autoencoder architecture (Ronneberger et al., 2015), to extract and normalize luminance and chrominance features, enabling robust SDR-to-HDR conversion. We specifically adopt the UNet architecture as the backbone of our encoder due to its ability to capture both global contextual information and fine-grained local details through its encoder-decoder design and skip connections. As pointed out by (Chen et al., 2021d), the SDR/HDR formation process can be effectively modeled as a combination of global and local operators, which aligns well with the inductive bias of UNet. In contrast, we experimented with a ResNet-based encoder and observed that it tends to introduce local structural artifacts and produce suboptimal gamut expansion, particularly in high-saturation regions. These limitations are primarily attributed to its lack of dedicated mechanisms for disentangling and recombining global and spatial context.
The encoder first extracts feature maps $e_{\text{lum}}$ and $e_{\text{chr}}$, which capture the luminance and chrominance characteristics of the input $x$, respectively. These feature maps are then passed through their corresponding projection heads, $\phi_{\text{lum}}$ and $\phi_{\text{chr}}$, to obtain normalized global and local representations: $z_{\text{lum}}^{\text{g}}$, $z_{\text{chr}}^{\text{g}}$, $z_{\text{lum}}^{\text{l}}$, and $z_{\text{chr}}^{\text{l}}$. The representation generation is formulated as:
(2) $z=\frac{\phi(F(x))}{\|\phi(F(x))\|_{2}},$
where $\phi(F(x))$ denotes the transformation of the feature map $F(x)$ via the corresponding projection head.
To learn representations that are aware of both luminance and chrominance modalities, each view-specific encoder $F_{\text{lum}}$ and $F_{\text{chr}}$ is paired with a structurally identical negative encoder, which encodes negative samples for contrastive learning. Meanwhile, the original feature maps $e_{\text{lum}}$ and $e_{\text{chr}}$ are preserved for the subsequent SDR-to-HDR transfer stage, as they retain essential spatial and semantic information required for high-quality reconstruction. These features are fed into a multi-view fusion module, which focuses on fusing the luminance and chrominance cues. The fusion process is refined using multi-head convolutions and incorporates a Sigmoid-based gating mechanism to dynamically balance the contributions of each component.
Algorithm 1 Luma-/Chroma-aware Negative Exemplars Generation
Input: Multi-degradation dataset $D$, parameters $k_{l}$ (luminance negatives), $k_{c}$ (chrominance negatives)
Output: Luminance negatives $N_{l}$, Chrominance negatives $N_{c}$
1:  Initialize $N_{l}\leftarrow[]$, $N_{c}\leftarrow[]$
2:  for each image $I\in D$ do
3:     Convert to YCbCr: Extract brightness ($L_{\text{ref}}$) and color components ($C_{\text{ref}}$)
4:     Initialize lists $List_{l}\leftarrow[]$, $List_{c}\leftarrow[]$
5:     for each degradation $x_{d}\in I.\text{degradations}$ do
6:        Extract brightness ($L_{d}$) and color components ($C_{d}$) of $x_{d}$
7:        Compute differences: $D_{l}\leftarrow\|L_{d}-L_{\text{ref}}\|_{1}$, $D_{c}\leftarrow\|C_{d}-C_{\text{ref}}\|_{1}$
8:        Append $(D_{l},x_{d})$ to $List_{l}$, $(D_{c},x_{d})$ to $List_{c}$
9:     end for
10:     Select top $k_{l}$ luminance-based and top $k_{c}$ chrominance-based samples
11:     $I_{l}\leftarrow\text{TopK}(List_{l},k_{l})$, $I_{c}\leftarrow\text{TopK}(List_{c},k_{c})$
12:     for each $(x_{l},x_{c})$ in $I_{l}$ and $I_{c}$ do
13:        Add $(L_{l},C_{\text{ref}})$ to $N_{l}$
14:        Add $(L_{\text{ref}},C_{c})$ to $N_{c}$
15:     end for
16:  end for
17:  Convert $N_{l}$, $N_{c}$ back to RGB
18:  Return $N_{l}$, $N_{c}$
3.1.2. Style Disentangled Representation Learning
The distributions of luminance and chrominance across various SDR degradations exhibit clear inconsistencies, as shown in Figure 4. Such inconsistencies can introduce undesirable interference into the representation space, hindering the model’s ability to learn robust features. To address this issue, we propose a Luma-/Chroma-aware negative exemplar generation strategy. By explicitly separating luminance and chrominance components in the YCbCr color space, this approach eliminates the adverse effects of inconsistent degradation trends and enables independent learning of meaningful features. The detailed procedure of this strategy is illustrated in Algorithm 1. This mechanism ensures the construction of a robust representation space, enhancing the model’s ability to capture true scene characteristics for effective SDR-to-HDR conversion.
Positive exemplars are generated by applying random horizontal and vertical flips to the anchor image. This augmentation strategy introduces spatial misalignment while preserving similar brightness and color characteristics, making the resulting samples suitable as positives for contrastive learning. The negative exemplars are designed to emphasize specific differences in brightness and color rather than directly using SDR degradations. Using the L1 norm, the top $k$ luminance-based and chrominance-based samples are selected as negative exemplars. To isolate these differences, additional samples are generated by replacing either the chrominance or luminance channels with those of the anchor. This approach ensures independent brightness and color variations, facilitating robust and disentangled representation learning.
Specifically, the encoder first extracts feature maps $e_{\text{lum}}$ and $e_{\text{chr}}$, representing the luminance and chrominance characteristics of the input image $I$, respectively. These features are then passed through corresponding non-linear projection heads $\phi_{\text{lum}}$ and $\phi_{\text{chr}}$, designed following the MoCo v2  (Chen et al., 2020a), to obtain the projected representations:
(3) $z_{\text{lum}}=\phi_{\text{lum}}(e_{\text{lum}}),\quad z_{\text{chr}}=\phi_{\text{chr}}(e_{\text{chr}}).$
These projected embeddings $z_{\text{lum}}$ and $z_{\text{chr}}$ serve as the basis for contrastive learning, enabling the model to construct a degradation-invariant and semantically meaningful feature space.
For negative samples $\{N_{\text{lum}}\}$ and $\{N_{\text{chr}}\}$, we define the proposed contrastive learning loss as follows:
(4) $\mathcal{L}_{\text{contra}}=\sum_{c\in\{\text{lum},\text{chr}\}}\sum_{s\in\{\text{g},\text{l}\}}-\log\left(\frac{e^{z_{c}^{s}\cdot z_{c}^{s+}}}{e^{z_{c}^{s}\cdot z_{c}^{s+}}+\sum_{j=1}^{N_{c}}e^{z_{c}^{s}\cdot z_{c,j}^{s}}}\right)$
where $c\in\{\text{lum},\text{chr}\}$ denotes the feature type, i.e., luminance (lum) or chrominance (chr); $s\in\{\text{g},\text{l}\}$ denotes the feature scope, i.e., global (g) or local (l); $z_{c}^{s}$ is the feature vector of the anchor sample for feature type $c$ and scope $s$; $z_{c}^{s+}$ is the positive sample, which is generated by applying data augmentation to the anchor sample $z_{c}^{s}$; $z_{c,j}^{s}$ are negative samples, where $j=1,2,\dots,N_{c}$, and $N_{c}$ is the total number of negative samples for feature type $c$.
By pulling the anchor sample and positive sample closer in the representation space while pushing the anchor sample and negative samples apart, the proposed method focuses on learning representations of the inherent luminance and chrominance characteristics of the scene, free from interference by specific styles. This process ensures that the model captures the true priors of luminance and chrominance distributions, effectively disentangling them from the variations introduced by different SDR styles. As a result, the learned representations reflect the underlying real-world luminance and chrominance properties, providing a robust foundation for achieving accurate HDR reconstruction.
3.2. Degradation-Domain Aware Controlled Mapping Network
Figure 5. Illustration of (a) Dense Controlled Mapping (DCM) and (b) Sparse Controlled Mapping (SCM).
3.2.1. Degradation-Aware Controlled Mapping
To address complex and spatially diverse degradations in SDR inputs, we introduce two hierarchical modulation modules: Dense Controlled Mapping (DCM) and Sparse Controlled Mapping (SCM). Both perform degradation-aware affine modulation on layer-normalized features, conditioned on global or local degradation priors. To ensure stable learning, we insert a zero-initialized convolution—termed the control layer—before injecting priors. Inspired by the soft control mechanism in ControlNet (Zhang et al., 2023), this layer acts as a learnable gate that gradually increases the influence of priors during training.
Directly injecting strong priors (e.g., contrastively pre-trained representations) early in the network often causes misalignment with low-level features, leading to instability. The control layer mitigates this by initially suppressing prior impact, allowing the network to first establish a stable restoration baseline. Over time, it adaptively integrates useful prior cues, improving both training stability and generalization to unseen or compound degradations.
Dense Controlled Mapping (DCM) globally modulates features based on a compact degradation representation. Given a global degradation code $z_{\text{deg}}^{g}\in\mathbb{R}^{C\times 1\times 1}$, we use two separate zero-initialized $1\times 1$ convolutions to generate the affine parameters $\gamma$ and $\beta$, i.e., $\gamma=\text{Conv}_{\gamma}^{1\times 1}(z_{\text{deg}}^{g})$, $\beta=\text{Conv}_{\beta}^{1\times 1}(z_{\text{deg}}^{g})$. These parameters modulate the normalized input feature via an affine transformation $\hat{x}=\gamma\cdot x+\beta$. The modulated feature is then refined by a $1\times 1$ convolution and added back to the input through a residual connection. The use of zero-initialized convolutions enables degradation priors to be injected in a stable and learnable manner over the training process. To further enhance normalization under diverse degradations, we extend this design to a degradation-aware Layer Normalization, where the affine parameters $\gamma$ and $\beta$ are dynamically generated from $z_{\text{deg}}^{g}$, allowing the network to adapt its normalization behavior based on the degradation type.
Sparse Controlled Mapping (SCM) handles spatially-varying degradations using a local degradation map $z_{\text{deg}}^{l}\in\mathbb{R}^{C\times H\times W}$. To generate pixel-wise affine parameters, we apply two parallel branches: each starts with a shared $1\times 1$ convolution followed by ReLU activation, and then a zero-initialized $3\times 3$ convolution to produce either $\gamma$ or $\beta$, i.e., $\gamma=\text{ZeroConv}_{\gamma}^{3\times 3}(\text{ReLU}(\text{Conv}^{1\times 1}(z_{\text{deg}}^{l})))$, and similarly for $\beta$. These parameters are used to modulate the normalized feature spatially as $\hat{x}=\gamma\odot x+\beta$, followed by a $1\times 1$ convolution and residual connection. This design allows fine-grained control of feature adaptation under heterogeneous degradations.
3.2.2. Detail Refinement
Although degradation-aware controlled mapping improves HDR reconstruction, residual discrepancies and high-frequency artifacts may still persist. To address these, we incorporate a Detail Refinement module as the final stage of the DDACMNet. Focused on enhancing texture and detail reconstruction, the refinement module employs Residual Blocks (ResBlocks) and is defined as:
(5) $F_{\text{refine}}(x)=\text{Conv}_{3\times 3}\circ\left[\text{ResBlock}\right]^{N_{r}}\circ\text{Conv}_{3\times 3}$
(6) $I_{\text{HDR}}=F_{\text{refine}}(I_{\text{LM}})$
where $N_{r}$ represents the number of stacked ResBlocks. This stage refines the output from the local mapping module ($I_{\text{LM}}$) to produce the final HDR image ($I_{\text{HDR}}$). The DR module ensures high-quality SDR-to-HDR conversion by reconstructing textures and resolving fine details left unaddressed by earlier stages.
Methods 2390eetf_gm 2446a 2446c_gm davinci hc_gm ocio2 reinhard youtube Average
[…]
HDRUNet 29.91 / 0.9219 25.55 / 0.8861 24.17 / 0.8509 23.17 / 0.8790 24.67 / 0.8548 25.12 / 0.8676 23.66 / 0.8887 22.76 / 0.8758 24.88 / 0.8781
HDRTVNet 28.97 / 0.9059 23.73 / 0.8502 26.05 / 0.8597 25.50 / 0.8815 26.91 / 0.8621 26.46 / 0.8699 23.23 / 0.8722 25.31 / 0.8709 25.77 / 0.8716
FMNet 27.96 / 0.9050 23.84 / 0.8753 22.29 / 0.8327 24.48 / 0.8835 23.16 / 0.8390 24.88 / 0.8603 22.26 / 0.8603 24.31 / 0.8755 24.15 / 0.8665
ICTCPNet 29.58 / 0.8548 27.02 / 0.8548 27.76 / 0.8594 25.67 / 0.8705 28.27 / 0.8612 29.97 / 0.8709 25.63 / 0.8694 25.71 / 0.8639 27.45 / 0.8631
LSNet 33.23 / 0.9408 31.18 / 0.9080 28.05 / 0.8731 25.63 / 0.9000 27.34 / 0.8712 29.56 / 0.8913 25.67 / 0.8970 27.01 / 0.9019 28.46 / 0.8979
AirNet 23.00 / 0.8594 21.72 / 0.8519 21.02 / 0.8044 22.22 / 0.8672 21.59 / 0.8107 22.18 / 0.8407 18.72 / 0.8379 22.20 / 0.8512 21.58 / 0.8404
PromptIR 30.06 / 0.9148 30.72 / 0.9073 28.23 / 0.8691 27.83 / 0.9028 28.00 / 0.8740 29.30 / 0.8838 24.97 / 0.8931 27.62 / 0.9068 28.34 / 0.8940
RAM-PromptIR 30.25 / 0.9188 30.83 / 0.9103 28.20 / 0.8711 28.07 / 0.9088 28.09 / 0.8770 29.31 / 0.8828 25.29 / 0.9001 27.77 / 0.9108 28.48 / 0.8975
Ours 34.13 / 0.9507 34.41 / 0.9333 30.38 / 0.8936 30.05 / 0.9266 30.47 / 0.8976 31.36 / 0.9083 27.59 / 0.9292 30.03 / 0.9358 31.05 / 0.9219
Table 1. Comparison of quantitative results on HDRTV4K (known degradation benchmark). All methods were trained using all degraded SDR versions from the HDRTV4K dataset. Bold and underline indicate the best and second-best results, respectively.
Metrics HDRUNet HDRTVNet FMNet ICTCPNet
PSNR 20.83 23.78 21.61 21.71
SSIM 0.8667 0.8824 0.8609 0.7961
Metrics LSNet PromptIR RAM-PromptIR Ours
PSNR 20.36 21.24 21.39 29.12
SSIM 0.8603 0.7946 0.8089 0.9412
Table 2. Comparison of quantitative results on HDRTV1K (unknown degradation benchmark). All methods were trained using all degraded SDR versions from the HDRTV4K dataset. Bold and underline indicate the best and second-best results, respectively.
3.3. Training Strategy
Our training strategy is tailored for the multi-view degradation encoder and the Degradation-Domain Aware Controlled Mapping network. Inspired by MoCo (He et al., 2019), the encoder is paired with a negative network whose weights are updated using Exponential Moving Average (EMA). This negative encoder helps maintain a slowly evolving representation space, providing stable and consistent targets for contrastive learning. The training process is divided into two stages, with specific adjustments in each stage to optimize learning.
In the first stage, contrastive learning is disabled, and the negative networks in the encoder are frozen to avoid premature noise injection. The control convolutions in the modulation modules are initialized as standard convolutions, and modulation is fully enabled. This allows the encoder and backbone to learn stable feature representations and basic degradation-aware modulation, serving as a strong initialization for the second stage. In the second stage, contrastive learning is introduced to improve the encoder’s discrimination of degradation patterns. The control convolutions are re-initialized as zero-initialized layers (ZeroConv), and gradually trained from zero to softly inject degradation priors. This progressive activation stabilizes the training and prevents disruption from strong contrastive priors, enabling effective adaptation to diverse degradations.
The pixel loss function is defined as:
(7) $L_{\text{pix}}=\|I_{\text{HDR}}^{\text{pred}}-I_{\text{HDR}}^{\text{G.T.}}\|_{1}+\lambda_{l}\|I_{\text{GM}}-I_{\text{HDR}}^{\text{G.T.}}\|_{1},$
where $I_{\text{HDR}}^{\text{pred}}$ is the predicted HDR image, $I_{\text{HDR}}^{\text{G.T.}}$ is the ground truth HDR image, and $\lambda_{l}$ is a hyper-parameter that balances the contributions of the global mapping (GM) loss and the final HDR prediction loss. Here, $I_{\text{GM}}$ denotes the intermediate output from the multi-layer DCM module in DDACMNet, which captures the globally modulated representation before detail refinement. Empirically, $\lambda_{l}$ is set to 0.7.
The total loss function integrates the pixel loss and contrastive loss as:
(8) $L_{\text{total}}=L_{\text{pix}}+\lambda_{\text{contra}}L_{\text{contra}}$
where $L_{\text{contra}}$ represents the contrastive loss for robust multi-view representation learning, and $\lambda_{\text{contra}}$ is empirically set to 0.2 to balance the contributions of the two loss terms.
Figure 6. Comparison of t-SNE visualizations for PromptIR (left) and our method (right). Each 2D coordinate represents the projection of a single frame’s feature vector from the corresponding dataset, with PCA (16 components) followed by t-SNE (perplexity=50, n_components=2). The visualizations highlight the structural differences between datasets, with the red points representing features from unknown degradations. This demonstrates our method’s ability to produce a more compact and distinct feature space, as shown by the separation and clustering of both known and unknown degradation types.
4. Experiments
4.1. Experimental Setup
4.1.1. Dataset.
We evaluate the effectiveness of our proposed framework on the HDRTV4K dataset (Guo et al., 2023), which consists of 3878 4K HDR images encoded in BT.2020/PQ1000. All degradation types included in HDRTV4K are utilized for training and testing. The training set contains 3478 images from HDRTV4K, while the testing set comprises 400 images from HDRTV4K. Additionally, we include 117 images from the HDRTV1K dataset (Chen et al., 2021d) in the testing set to evaluate the performance on unknown degradations, resulting in a total of 517 test images.
4.1.2. Implementation Details.
Each multi-view encoder outputs a global vector and a local feature map with 64 and 16 channels, respectively. The mapping network consists of convolutional layers with 32 channels. We implement our framework using the PyTorch library and adopt the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a learning rate initialized to $2\times 10^{-4}$. The learning rate is scheduled using a MultiStepLR strategy, with decay milestones at 200K and 400K iterations and a decay factor of 0.5. We employ exponential moving average (EMA) with a decay factor of 0.999 to stabilize training. The model is trained for 400K iterations in total, with the first 40K iterations designated as Stage 1. Mixed-precision training with bfloat16 is enabled to improve computational efficiency. We use a mini-batch size of 16, and the training is conducted on an NVIDIA RTX 4090 GPU, taking approximately 3 days to complete.
4.2. Experimental Results
Figure 7. Visual comparison with the state-of-the-art methods. Following hdrtvdm (Guo et al., 2023), the HDR images are shown in their original encoding without tone mapping to preserve highlights. Transfer function and gamut mismatches may cause darker, less saturated grayish appearance on paper or SDR screens.
4.2.1. Quantitative Results.
We validate RealRep on the HDRTV4K (known degradation benchmark) and HDRTV1K (unknown degradation benchmark) datasets, comparing it to state-of-the-art SDR-to-HDR methods, including HDRUNet(Chen et al., 2021a), HDRTVNet(Chen et al., 2021d), FMNet(Xu et al., 2022), ICTCPNet(Huang et al., 2023), LSNet(Guo et al., 2023), PromptIR(Potlapalli et al., 2023), and RAM-PromptIR(Qin et al., 2024).
As shown in Table 1, RealRep outperforms all competing methods on HDRTV4K, demonstrating its ability to handle diverse degradations and produce high-quality HDR reconstructions. This stems from its disentangled multi-view representation learning, which separates luminance and chrominance distributions, enabling robust adaptation to varying degradation characteristics. On HDRTV1K (Table 2), RealRep excels in generalizing to unknown degradations, outperforming methods like HDRUNet, HDRTVNet, and FMNet, which suffer significant performance drops. PromptIR, while effective in simpler tasks like noise removal, struggles with complex luminance and chrominance transformations. RealRep consistently delivers superior HDR outputs, leveraging its disentangled encoder and domain-adaptive transfer framework.
4.2.2. Qualitative Results.
Figure 7 compares HDR/WCG images generated by state-of-the-art methods and our proposed RealRep framework. Existing methods, such as HDRTVNet, FMNet, and ICTCPNet, struggle with diverse degradations, causing artifacts like banding, color distortion, and uneven luminance due to their reliance on fixed tone mapping strategies, which limits adaptability and leads to inconsistent performance. For instance, HDRTVNet produces overexposure and banding, FMNet shows luminance and chrominance mismatches, and ICTCPNet renders desaturated colors. In contrast, RealRep combines disentangled multi-view degradation representation learning with a domain-adaptive hierarchical transfer framework to dynamically adapt to varying degradations, producing HDR images that are natural, consistent, and artifact-free. These qualities are evident in Figure 6, where RealRep achieves smoother luminance transitions, better highlight recovery, and accurate color distributions, surpassing other methods in addressing unknown degradations and delivering results closer to the ground truth.
4.2.3. Joint Luma-/Chroma-aware Representation Visualization.
To demonstrate the effectiveness of RealRep in constructing the inverse tone mapping (iTM) embedding space and its generalization to unknown degradations, we performed t-SNE visualization of the feature space. As shown in Figure 3, the left panel depicts the prompt representations extracted by PromptIR, while the right panel illustrates the degradation features extracted by RealRep. In PromptIR’s feature space, unknown degradations (e.g., HDRTV1K) are scattered and overlap with other clusters, indicating poor generalization to unseen scenarios. In contrast, RealRep produces compact and well-separated clusters for HDRTV1K, showcasing superior generalization to unknown degradations and a more structured feature space. These findings highlight the strength of RealRep in addressing diverse real-world degradations.
Configuration 2390eetf_gm 2446a 2446c_gm davinci Avg.
Baseline 27.37 27.72 25.96 25.67 25.74
RealRep w/o fusion 32.27 32.76 29.68 27.71 29.83
RealRep w/o $z_{\text{lum}}$ 32.96 27.88 27.53 29.37 28.96
RealRep w/o $z_{\text{chr}}$ 32.92 29.48 28.40 29.42 29.58
RealRep w/o $\mathcal{L}_{\text{contra}}$ 33.73 29.71 28.23 29.77 30.04
RealRep 34.29 34.83 31.10 30.17 31.42
Table 3. Ablation on multi-view representations and multi-view fusion, evaluated by PSNR (dB). The final column reports the average across all eight degradation types. Due to space limitations, only four representative degradations are shown.
Configuration 2390eetf_gm 2446a 2446c_gm davinci hdrtv1k Avg.
RealRep w/o $z_{\text{global}}$ 24.56 23.43 28.41 22.41 25.03 25.00
RealRep w/o $z_{\text{local}}$ 32.95 34.00 30.90 27.33 28.42 30.28
RealRep w/o control 33.24 33.55 30.35 29.49 28.37 30.47
RealRep 34.29 34.83 31.10 30.17 29.12 31.16
Table 4. Ablation on the degradation-aware controlled mapping module, evaluated by PSNR (dB).
4.3. Ablation Studies.
In this section, we conduct a series of ablation experiments to investigate the effectiveness of the decoupled multi-view representations and degradation-aware controlled mapping module.
4.3.1. Ablation of Disentangled Multi-View Representations
To evaluate the effectiveness of disentangled multi-view representations, we conduct ablation experiments on the luminance representation $z_{\text{lum}}$, chrominance representation $z_{\text{chr}}$, their combination, the contrastive loss $\mathcal{L}_{\text{contra}}$, and the multi-view fusion module. As shown in Table 3, the baseline model, which lacks all multi-view representations and fusion, achieves the lowest average PSNR of 25.74dB, highlighting the necessity of these components. Removing either $z_{\text{lum}}$ or $z_{\text{chr}}$ reduces the average PSNR to 28.96dB and 29.58dB, respectively, underscoring their independent contributions. The absence of the multi-view fusion module, which integrates luminance and chrominance features, further lowers the average PSNR to 29.83dB, demonstrating its importance for balancing and integrating these representations. Furthermore, removing the contrastive loss $\mathcal{L}_{\text{contra}}$ leads to a performance drop from 31.42dB to 30.04dB, suggesting that the contrastive learning space plays a crucial role in guiding the model to extract informative and discriminative priors. The full RealRep model achieves the highest average PSNR of 31.42dB, validating the joint effectiveness of multi-view representations, fusion, and contrastive supervision in enhancing SDR-to-HDR performance.
4.3.2. Ablation of Degradation-Domain Aware Controlled Mapping Network
We further investigate the impact of the degradation-aware controlled module by conducting ablation experiments on its global and local mapping components, as well as the control mechanism. As shown in Table 4, removing either the global mapping or local mapping leads to a noticeable performance drop, with average PSNRs of 25.00dB and 30.28dB, respectively, validating their complementary roles in capturing hierarchical domain knowledge. Notably, the control mechanism—comprising a zero-initialized convolution and a two-stage training strategy—plays a key role in enhancing generalization. Removing it leads to a decline in both overall performance and generalization ability, with the average PSNR dropping from 31.16dB to 30.78dB, and the PSNR on the challenging HDRTV1K dataset decreasing from 29.12dB to 28.87dB. These results highlight the effectiveness of the control design in improving transferability and robustness across diverse and unseen degradation scenarios. The full RealRep model consistently achieves the best results, demonstrating the importance of the complete degradation-aware controlled mapping design.
4.3.3. Ablation of encoder architecture
We compare ResNet and our UNet-based encoder under identical training settings (batch size = 16, 100K iterations). As shown in Table 5, the UNet encoder achieves better performance (27.28 dB vs. 26.87 dB) with fewer parameters (0.50M vs. 1.23M), demonstrating its superior efficiency and representation capability.
Encoder Parameters (M) PSNR (dB)
ResNet Encoder 1.2289 26.87
UNet Encoder (Ours) 0.5044 27.28
Table 5. Ablation on encoder architecture. UNet-based encoder outperforms ResNet under identical training settings.
5. Concluding Remarks
This paper presents RealRep, a novel SDR-to-HDR method designed for real-world scenarios with diverse SDR styles. RealRep tackles luminance and chrominance variations using style disentangled representation learning, capturing intrinsic style differences to provide realistic guidance for inverse tone mapping. A Degradation-Domain Aware Controlled Mapping framework further enables dynamic luminance range expansion and color gamut extension. Evaluations on diverse and unknown degradations demonstrate our superiority over state-of-the-art methods in visual quality and generalization. Future work will explore extending RealRep to video HDR conversion and real-time applications.